OpenStudy (anonymous):

When the quadratic equation y = (x − 3)(4x + 2) is written in standard form, what is the value of the coefficient "b?"

OpenStudy (anonymous):

y = (x − 3)(4x + 2) y = 4x² + 2x − 12x − 6 y = 4x² − 10x − 6 y = ax² + bx + c ⇒ b = −10 <<<------Answer

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

i would have never got that

OpenStudy (anonymous):

Yes it was easy just take your time and do the equation on paper than redo it the other way to double check your answer.